OpenStudy (anonymous):

Find the range of f(x) = -x+4 for the domain (-3,-2,-1,1)

OpenStudy (anonymous):

plug in these x values into your equation, range goes from lowest value to highest value

OpenStudy (anonymous):

f(-3)=-3+4 f(-3)=1 like that?

jimthompson5910 (jim_thompson5910):

Close, more like f(x) = -x+4 f(-3) = -(-3)+4 f(-3) = 3+4 f(-3) = 7

jimthompson5910 (jim_thompson5910):

Do this for all the given x values. The corresponding y values or f(x) values will form a new set which will be the range.

OpenStudy (anonymous):

I think i have it now. thank you (:

jimthompson5910 (jim_thompson5910):

sure thing